OpenStudy (anonymous):

Easy Quotient/Product Rule Question f(x) = (x^4)(cos x)^-1 I know the answer is x^4 tan x + 4x^3 sec x But when I did the problem again today, I kept coming up with a different answer. Could someone solve with steps?

OpenStudy (campbell_st):

using the product rule u = x^4 du/dx = 4x^3 v = (cos(x))^-1 dv/du = -1 (cos(x))^-2 * -sin(x) so \[\frac{dy}{dx} = u \times \frac{dv}{dx} + v \times \frac{du}{dx}\] which gives \[\frac{dy}{dx} = x^4 \times \sin(x)(\cos(x)^{-2} + (\cos(x))^{-1} \times 4x^3\] which can be simplified to \[\frac{dy}{dx} = \frac{x^4 \sin(x)}{(\cos(x))^2} + \frac{4x^3}{\cos(x)}\]

OpenStudy (anonymous):

Thanks so much!